A spring 20cm long is stretched to 25cm by a load of 50N. What will be its length when stretched by 100N assuming that the elastic limit is not reached


1
Expert's answer
2021-03-08T08:11:43-0500

"\Delta x=0.25-0.20=0.05\ (m)"


"F=k\Delta x\to k=F/\Delta x=50/0.05=1000\ (N/m)"


"\Delta x_1=F_1/k=100/1000=0.1\ (m)"


"x_1=0.2+0.1=0.3\ (m)" . Answer